OpenStudy (anonymous):

The nitrates of beryllium, calcium, magnesium, and strontium all decompose in the same way when heated. When 2.00 g of one of these anhydrous nitrates is decomposed, 1.32 g of gas is produced. What is the nitrate? A beryllium nitrate B calcium nitrate C magnesium nitrate D strontium nitrate

thomaster (thomaster):

These are all group 2 elements. Their nitrates will decompose to the oxide, nitrogen dioxide and oxygen This will happen: (replace the X with Be, Ca, Mg, Sr) \(\large\sf 2X(NO_3)_2(s)\to 2XO(s)+4NO_2(g)+O_2(g)\) So for beryllium: \(\to\sf 2Be(NO_3)_2 ~\to 2BeO ~+ 4NO_2 + O_2\) Calcium: \(\to\sf 2Ca(NO_3)_2 ~\to 2CaO ~+ 4NO_2 + O_2\) Magnesium: \(\to\sf 2Mg(NO_3)_2 \to 2MgO + 4NO_2 + O_2\) Strontium: \(\to\sf 2Sr(NO_3)_2 ~~\to 2SrO ~~+ 4NO_2 + O_2\) I'll give you an example (it's not the correct answer but a way you can get to the answer) Magnesiumnitrate has a molecular weight of 148.30 g/mol. You can use the ratio to calculate the amount of gas produced. 2 moles of \(\sf 2Mg(NO_3)_2\) will produce 4 moles \(\sf NO_2\) and 1 mole \(\sf O_2\). So you have 2 grams of berylliumnitrate. First thing you do is calculate how many moles that is. To do that you divide by molecular weight. So \(\sf \frac{2}{148.3}\) =0.01348 moles of magnesiumnitrate. So the ratio is 2 \(\to\) 4 :1 when 2 = 0.01348, you double that to get the 4, so you have 0.02697 mol \(\sf NO_2\). To get the 1 you halve it. so 0.0067425g \(\sf O_2\) The molecular weight of \(\sf NO_2\) = 46 g/mol MW of \(\sf O_2\) is 32 g/mool To get the grams you multiply with the amount of moles. So for \(\sf NO_2 \to\) 0.02697*46= 1.24062 gram nitrogen dioxide for \(\sf O_2\) 0.0067425*32 = 0.21576 gram oxygen Together thats 1.24062+0.21576 = 1.45638 gram gas So now you know that C is not the answer. Do this for A, B & D and you'll find out the correct answer (you can find molecular weights on http://ptable.com) \(\overline{\underline{\LARGE{\color{gold}{\star~}}\Large\tt\color{green}{I\;Hope\;this\;Helps!}\LARGE{\color{gold}{~\star}}}}\)

thomaster (thomaster):

link: http://ptable.com (other one has a parenthesis included :P)

thomaster (thomaster):

And sorry i said berylliumnitrate in the 5th part. supposed to be magnesiumnitrate \(\Huge\ddot\smile\)

OpenStudy (anonymous):

thankyou :)